OpenStudy (ccastudentgrade10):

help please asap will medal/fan!! solve using elimination 5x + 4y = 12 3x - 3y = 18 please explain/show each step @texaschic101 @vocaloid

OpenStudy (jango_in_dtown):

Multiply the first equation by3 and the second equation by 4 and add them

OpenStudy (jango_in_dtown):

you will get the value of x . and hence putting the value of x in any of the equations you will obtain y

OpenStudy (texaschic101):

5x + 4y = 12 ---> multiply by 3 3x - 3y = 18 ---> multiply by 4 ---------------- 15x + 12y = 36 (result of multiplying by 3) 12x - 12y = 72 (result of multiplying by 4) --------------add 27x = 108 x = 108/27 x = 4 now sub 4 in for x in either of beginning equations 5x + 4y = 12 5(4) + 4y = 12 20 + 4y = 12 4y = 12 - 20 4y = -8 y = -8/4 y = -2 check.. 3x - 3y = 18 3(4) - 3(-2) = 18 12 + 6 = 18 18 = 18 (correct) so x = 4 and y = -2 questions ?

OpenStudy (ccastudentgrade10):

wow that makes sense now, thank you @texaschic101, can i ask you one more question?